# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

### v1

@@ 5.1 Experimental Setup
We train URS on $11$ mixed VRP variants to ensure that each feature in UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. Specifically, training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024b); (9) OCVRP (Zhou et al., 2024b); (10) CVRPB (Zhou et al., 2024b), and (11) OCVRPTW (Zhou et al., 2024b). To evaluate the zero-shot generalization of URS, we construct 96 unseen variants to further highlight the broad applicability of URS. And each test dataset consists of 1000 instances. The oracle solver for each problem can be found in Appendix D.
@@
We use an embedding dimension $d$ of $128$ and a feed-forward layer dimension of $512$. We set the number of attention layers in the encoder to $12$. The clipping parameter $\zeta=50$ in Equation 9 for better training convergence (Jin et al., 2023). For training, URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details). We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and the batch size is $128$. At each training step, we randomly select one problem $P_{i}$ from the problem set $\mathcal{P}$. We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch. Weight decay is set to $10^{-6}$. More details about the model training are provided in Appendix H.
@@
We compare URS with the following methods: (1)Classical Solvers: [^1] We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019). For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a). PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron & Furnon, 2023); (2)Neural Solvers: (i) specialized solvers that support multiple VRP variants: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2024a) and POMO (Kwon et al., 2020); (ii) general neural solvers capable of handling multiple variants [^2] Given the differences in problem variants and visible task training among the baseline models, we only compare methods with the same training setup. Thus, RF Berto et al. (2024) and CaDA (Li et al., 2025a) are not included in our experiments to ensure a fair comparison.: MTPOMO (Liu et al., 2024a), MVMoE (Zhou et al., 2024b), ReLD-MTL (Huang et al., 2025), GOAL-MTL (Drakulic et al., 2025); (iii) Solvers for less-explored variants: MatNet (Kwon et al., 2021), Heter-AM (Li et al., 2021).
@@
We report the average objective value (Obj.), optimality gap (Gap), and total inference time (Time) for each method. The optimality gap quantifies the discrepancy between the solutions generated by the corresponding methods and the near-optimal solutions, which are obtained using classical solvers. For most NCO baseline methods, we execute the source code provided by the authors using default settings. We compare the performance of URS to the relevant baselines for each problem as well as its single-task version (i.e., URS-STL). For symmetric instances, we report the best result with $\times 8$ instance augmentation, following Kwon et al. (2020). For asymmetric instances, we report the best result with $\times 128$ instance augmentation, following Kwon et al. (2021).
Table 2: Performance on 1K test instances of $11$ seen VRPs.
Method ATSP TSP OP PCTSP PDTSP -
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 2m 0.00% 6m 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m -
Sym-NCO - 0.14% 5s 0.68% 15s 0.14% 16s - -
BQ 1.27% 1m 0.35% 7s 0.63% 7s - - -
LEHD - 0.59% 2s - - - -
POMO - 0.13% 5s - - - -
ICAM 3.05% 1m 0.15% 4s - - - -
MatNet 0.95% 6.5m - - - - -
Heter-AM - - - - 6.61% 5m -
GOAL-MTL 1.77% 1m - 1.20% 38s - - -
[…]
URS 2.26% 1.1m 0.57% 6s 0.45% 4s 1.06% 5s 4.98% 4s -
Method ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 9.1m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
Sym-NCO - 1.46% 6s - - - -
BQ - 3.24% 8s - - - -
LEHD - 3.68% 2s - - - -
[…]
ICAM - 2.04% 4s - - - -
MVMoE - 1.65% 12s 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
MTPOMO - 1.85% 6s 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
[…]
GOAL-MTL - 4.22% 48s 4.66% 42s - - -
[…]
URS 3.06% 1.4m 1.81% 6s 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
Table 3: Zero-shot generalization on 1K test instances of $30$ unseen VRPs. Due to page limit, complete experimental results (including all 96 unseen variants) are deferred to Appendix I.
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
[…]
Method MDCVRP MDCVRPL MDOCVRPB MDOCVRPBPTW MDOCVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 31.90% 31s 33.42% 44s 28.21% 26s 46.18% 40s 45.23% 40s
MTPOMO 23.75% 27s 24.32% 25s 25.06% 20s 34.19% 29s 33.37% 31s
ReLD-MTL 17.37% 36s 18.48% 28s 18.50% 21s 34.08% 33s 33.83% 36s
[…]
Method ACVRPB ACVRPL ACVRPTW AOPDCVRP ACVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBPTW ACVRPBPLTW ACVRPBPTW AOCVRPBPLTW ACVRPBTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDCVRPL AMDCVRPBPLTW AMDOCVRPBPLTW PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBPTW AMDCVRPBPTW AMDCVRP APDTSP APDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
@@ 5.2 Performance Evaluation
The experimental results on trained VRPs with scale $100$ are reported in Table 2. We observe that among all comparable neural solvers, URS-STL achieves the best performance on 8 out of the 11 routing problems, complemented by remarkably fast inference, highlighting the strength of the URS architecture when fully specialized. Meanwhile, mixing all problems into the training of URS incurs only a minor performance drop, yet still delivers competitive performance across all problems, demonstrating that potential connections among different problems are effectively captured by URS.
@@
We evaluate our URS model on zero-shot generalization across 96 unseen VRP variants, including 44 well-studied VRP variants from the NCO community and 52 newly introduced variants that have not been explored before. The detailed results are presented in Table 3 and Appendix I. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address previously unexplored variants while maintaining solution quality, which compounds asymmetry and PD relations. Without any fine-tuning, URS still constructs high-quality solutions for more than 90 VRP variants that are not seen in training, highlighting its strong zero-shot generalization ability.
@@
We further evaluate the generalization ability of URS on benchmark instances from CVRPLIB Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$, and we compare URS against several representative single-task and multi-task models. As demonstrated in Table 12 (Appendix J), URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based specialized model LEHD (8.678% vs. 12.836%), underscoring its practical applicability in real-world scenarios.
@@ 6 Ablation Study
Table 4: Effects of different components of URS.
Method Avg.gap
w/o $\bm{\xi}$ 3.78%
w/o $C_{t}$ 2.97%
$\{\bm{D}\}$ 6.33%
$\{\bm{D},\bm{D^{\mathrm{T}}}\}$ 4.07%
$\{\bm{D},\bm{R}\}$ 7.34%
POMO-Dec. 3.39%
ReLD-Dec. 3.87%
[…]
In this section, we conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures. As detailed in Table 4. URS maintains its competitive performance without $C_{t}$. And its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$. For MBM, the results reveal that three priors $\{\bm{D},\bm{D^{\mathrm{T}}},\bm{R}\}$ are complementary, and their combined use allows MBM to effectively learn the geometric and relational biases inherent in various problems, resulting in superior overall performance. Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while the integration of the decoder in ICAM (Zhou et al., 2024a) leads to a further improvement. More detailed results and analyses can be found in Appendix K.
@@ 7 Conclusion, Limitation, and Future Work
In this work, we propose a novel RL-based Unified Routing Solver (URS) for cross-problem zero-shot generalization. The core of URS is a Unified Data Representation (UDR), which broadens problem coverage by substituting problem enumeration with data unification. Its generalization capability is further enhanced by a Mixed Bias Module (MBM) and an adaptive parameter generator. Moreover, an LLM-based constraint satisfaction mechanism guarantees solution feasibility and reduces dependency on domain expertise. Extensive experiments demonstrate that URS can produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, more than 90 of which are unseen during training.
@@
While URS achieves impressive results on a wide range of VRP variants, its current training scheme does not accommodate their inherent differences in complexity. In the future, we aim to explore more effective training schemes to further improve training efficiency.
@@ A.1 Neural Routing Solvers with Single-Task Learning
The NCO methods have emerged as a promising paradigm for tackling various routing problems, achieving impressive empirical performance. Early seminal work based on Pointer Networks ignited interest in learning end-to-end construction policies for TSP and CVRP (Vinyals et al., 2015; Bello et al., 2016; Nazari et al., 2018). Subsequently, the Transformer-style heavy-encoder and light-decoder auto-regressive architecture has become the dominant paradigm (Kool et al., 2019), while POMO (Kwon et al., 2020) exploits permutation symmetry to enhance exploration diversity. Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a; b; Sun et al., 2024). However, models trained on a fixed instance scale often degrade sharply when evaluated out of scale. To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2024a), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025). In parallel, ”heavy decoder” designs that dynamically recompute embeddings at each construction step (Drakulic et al., 2023; Luo et al., 2023) demonstrate strong generalization on large-scale instances. To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024b; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun & Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023). Moreover, heterogeneous vehicle routing (Li et al., 2022b), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup–delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results. Despite these advancements, they still train a specialized model per problem type, limiting cross-problem transfer. We instead focus on a unified neural routing solver that aims to generalize across multiple routing problems without retraining from scratch for each specification.

### v2

@@ 4.1 Experimental Setup
We train URS on $11$ mixed variants to ensure that most features in UDR have been seen at least twice, which prevents URS from overfitting a specific feature to a single problem. Training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024); (9) OCVRP (Zhou et al., 2024); (10) CVRPB (Zhou et al., 2024), and (11) OCVRPTW (Zhou et al., 2024). We provide further discussion of problem selection in Appendix J.1. Each of 110 variants comprises 1,000 instances. The related oracle solvers and problem representations are detailed in Appendix D.
@@
We use an embedding dimension $d$ of $128$ and a feed-forward layer dimension of $512$. We set the number of attention layers in the encoder to $12$. The clipping parameter $\zeta$ is set to $50$ in Equation 10. URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details). We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and each batch has a size of $128$. At each training step, we randomly select one problem from $11$ problems. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch. Weight decay is set to $10^{-6}$. More details are provided in Appendix H.
@@
We compare URS with the following methods: (1)Classical Solvers: PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron and Furnon, 2023). We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019). For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a); (2)Neural Solvers: (i) specialist solvers: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2026), POMO (Kwon et al., 2020), ELG (Gao et al., 2024), MatNet (Kwon et al., 2021), and Heter-AM (Li et al., 2021); (ii) generalist neural solvers: MTPOMO (Liu et al., 2024), MVMoE (Zhou et al., 2024), ReLD-MTL (Huang et al., 2025), and GOAL-MTL (Drakulic et al., 2025). Given the disparities in problem formulations and training settings, detailed comparisons against RouteFinder (Berto et al., 2025) and CaDA (Li et al., 2025a) are deferred to Appendix J.2.
@@
We report the optimality gap (Gap) and total inference time (Time) for each method. The optimality gap quantifies the discrepancy between the corresponding solvers and oracle solvers. For neural solvers, we execute the source code provided by the authors using default settings. We compare the performance of URS with relevant baselines for each problem, as well as its single-task version (i.e., URS-STL). The URS-STL variant uses exactly the same architecture and hyperparameters (see Appendix H.2 for more details). For URS, we report the best result with $\times 8$ (Kwon et al., 2020) and $\times 128$ (Kwon et al., 2021) instance augmentation for symmetric and asymmetric instances, respectively. The mark (-) indicates that the method does not support the corresponding setting.
@@ 4.2 Performance Evaluation
As presented in Table 2 and Table 3, among comparable neural solvers, URS-STL achieves the lowest gap on 8 of the 11 VRPs while offering fast inference, highlighting the strength of the URS architecture when fully specialized. Although URS-MTL exhibits a performance degradation relative to STL variants, it remains competitive across all 11 problems. Using ATSP as an example, we attribute this drop to geometric structural inequality, which likely biases the model’s learning capacity toward symmetric VRPs. While the lack of explicit constraint states yields weaker performance compared to ReLD-MTL and MVMoE, URS shows clear advantages when generalizing to unseen VRP variants (Table 4) and scaling to large-size instances (Table 5).
@@
We evaluate our URS on zero-shot generalization across 99 unseen VRP variants. As presented in Table 4 and Appendix I, on 44 widely-studied VRPs with diverse constraint combinations (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given set of feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address 55 less-explored VRP variants that compound asymmetry and PD relations. These results highlight URS’s strong capacity for cross-problem zero-shot generalization without any fine-tuning.
Scalability to Larger-scale VRP Variants
[…]
@@
We further evaluate URS on benchmark instances from CVRPLib Set-X (Uchoa et al., 2017) and CVRPLib Set-AGS (also known as Set-XXL) (Arnold et al., 2019). We compare URS against several representative specialist and generalist solvers. As presented in Table 6, URS can significantly outperform existing generalist neural routing solvers in large-scale benchmark instances. Notably, as a generalist solver, URS also greatly surpasses many representative specialist models (e.g., LEHD, BQ, and ELG). In addition, the detailed results (Appendix J.3) on the Set-X dataset show that URS still maintains its position as the best overall performance. These results further underscore the practical applicability of URS in real-world large-scale scenarios.
Table 6: Comparison on Set-AGS with $N\in[3000,7000]$. All models are trained on instances of size $N=100$.
Type Method Leuven1 Leuven2 Antwerp1 Antwerp2 Avg.gap
(N=3000) (N=4000) (N=6000) (N=7000)
Specialist LEHD greedy 16.60% 34.86% 14.66% 22.77% 22.22%
BQ greedy 18.53% 30.70% 16.48% 27.67% 23.34%
[…]
GOAL greedy OOM OOM OOM OOM -
[…]
@@ 5 Ablation Study
In this section, we conduct a comprehensive ablation study and analysis to demonstrate the effectiveness and robustness of URS. The detailed results and discussion are presented below. Note that, unless otherwise stated, the instance augmentation technique is used for the ablation studies.
Effects of Key Components
We conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures. As detailed in Figure 4, URS maintains its competitive performance without $C_{t}$; its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$. For MBM, the results reveal that three priors $\{\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\}$ are complementary, and their combined use allows MBM to effectively learn the geometric and relational biases inherent in various problems, resulting in superior overall performance. Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while integrating the decoder introduced in Zhou et al. (2026) further improves cross-problem performance. More detailed results and analyses can be found in Appendix K.
Figure 4: Effects of different components of URS.
Table 7: Results comparing performance with and without MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ on seen and unseen variants.
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ Avg.gap Best
Seen (11) Unseen (11) Solution
$\times$ $\checkmark$ $\checkmark$ 6.33% 10.41% 3/22
$\checkmark$ $\times$ $\times$ 4.20% 8.12% 0/22
$\checkmark$ $\times$ $\checkmark$ 3.84% 8.02% 1/22
$\checkmark$ $\checkmark$ $\times$ 3.63% 7.81% 1/22
$\checkmark$ $\checkmark$ $\checkmark$ 2.74% 7.05% 17/22
MBM & Parameter Generator
To further evaluate the effects of MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ for cross-problem zero-shot generalization, we conduct tests on 11 seen problems and 11 unseen CVRP variants widely studied in (Liu et al., 2024). As shown in Table 7, the full URS model achieves the lowest average gap on both seen and unseen problems, and provides the best solution on the majority of tasks (17/22). This shows that their synergy is essential for robust cross-problem zero-shot generalization. We provide detailed results in Appendix K.5.
Table 8: Performance comparison of ablating each UDR component (i.e., $\bm{\rho}$, $\bm{\omega}$, and $\bm{\xi}$) on seen and unseen variants. Since removing positional identifiers $\bm{\rho}$ prevents the use of instance augmentation, this augmentation technique is disabled across all results to maintain a consistent inference strategy.
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ Avg.gap Best
Seen (11) Unseen (11) Solution
$\times$ $\checkmark$ $\checkmark$ 4.65% 18.12% 3/22
$\checkmark$ $\times$ $\checkmark$ 9.81% 18.60% 0/22
$\checkmark$ $\checkmark$ $\times$ 5.43% 18.60% 4/22
$\checkmark$ $\checkmark$ $\checkmark$ 4.17% 8.51% 15/22
Effects of Each UDR Component
To clarify the independent contribution of each UDR component, we conduct ablations by removing each component in turn. As shown in Table 8, the absence of any single component leads to notable performance degradation. A complete UDR achieves the lowest average gap, affirming the necessity of our UDR design for robust zero-shot generalization. The detailed results for each VRP variant can be found in Appendix K.6.
Hypernetwork vs. Adapter
We train URS-Adapter, a URS variant that adopts a distinct decoder for each of the 11 training problems, and compare it with the original URS with a hypernetwork on the same 11 seen problems and large-scale CVRP variants (see Appendix K.7 for detailed results). URS consistently outperforms URS-Adapter across the 11 problems seen, achieving higher parameter efficiency. This performance advantage holds consistently across large-scale CVRP variants, further solidifying our conclusion.
@@ 6 Conclusion, Limitation, and Future Work
In this work, we propose a novel RL-based URS for cross-problem zero-shot generalization. The core of URS is the proposed UDR, which broadens problem coverage by substituting problem enumeration with data unification. Its generalization capability is further enhanced by an MBM and a problem-conditioned parameter generator. Extensive experiments show that URS can produce high-quality solutions for 110 VRP variants without any fine-tuning, including 99 unseen variants, significantly broadening the problem coverage and reducing reliance on domain expertise.
@@
While URS achieves impressive results across a wide range of VRP variants, its current training scheme does not fully account for the inherent disparities in problem complexity. In the future, we aim to explore more effective strategies, such as hierarchical training and difficulty-adaptive sampling, to improve training efficiency. Additionally, extending our method to accommodate more complex variants(e.g., dynamic VRP) is a promising direction for future research.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62476118), the Guangdong Provincial Key Laboratory of Fully Actuated System Control Theory and Technology (Grant No. 2024B1212010002), the Natural Science Foundation of Guangdong Province (Grant No. 2024A1515011759), and the Center for Computational Science and Engineering at Southern University of Science and Technology. We would like to thank the anonymous reviewers and (S)ACs of ICML 2026 for their constructive comments and dedicated service to the community. Additionally, we would like to thank Mr. Rongsheng Chen for his valuable advice and assistance during this research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Neural Routing Solvers with Single-Task Learning
The NCO methods have emerged as a promising paradigm for tackling various routing problems, achieving impressive empirical performance. Early seminal work based on Pointer Networks ignited interest in learning end-to-end construction policies for TSP and CVRP (Vinyals et al., 2015; Bello et al., 2016; Nazari et al., 2018). Subsequently, the Transformer-style heavy-encoder and light-decoder auto-regressive architecture has become the dominant paradigm (Kool et al., 2019), while POMO (Kwon et al., 2020) exploits permutation symmetry to enhance exploration diversity. Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a, b; Sun et al., 2024). However, models trained on a fixed instance scale often degrade sharply when evaluated out of scale. To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2026), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025). In parallel, ”heavy decoder” designs that dynamically recompute embeddings at each construction step (Drakulic et al., 2023; Luo et al., 2023) demonstrate strong generalization on large-scale instances. To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun and Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023). Moreover, heterogeneous vehicle routing (Li et al., 2022), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup-delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results. Despite these advancements, they still train a specialist model per problem type, limiting cross-problem transfer. We instead focus on a unified neural routing solver that aims to generalize across multiple routing problems without retraining from scratch for each specification.